### v2

# IMU: Influence-guided Machine Unlearning

Abstract
Recent studies have shown that deep learning models are vulnerable to attacks and tend to memorize training data points, raising significant concerns about privacy leakage. This motivates the development of machine unlearning (MU), i.e., a paradigm that enables models to selectively forget specific data points upon request. However, most existing MU algorithms require partial or full fine-tuning on the retain set. This necessitates continued access to the original training data, which is often impractical due to privacy concerns and storage constraints. A few retain-data-free MU methods have been proposed, but some rely on access to auxiliary data and precomputed statistics of the retain set, while others scale poorly when forgetting larger portions of data. In this paper, we propose Influence-guided Machine Unlearning (IMU), a simple yet effective method that conducts MU using only the forget set. Specifically, IMU employs gradient ascent and innovatively introduces dynamic allocation of unlearning intensities across different data points based on their influences. This adaptive strategy significantly enhances unlearning effectiveness while maintaining model utility. Results across vision and language tasks demonstrate that IMU consistently outperforms existing retain-data-free MU methods.
1 Introduction
Deep learning (DL) models are widely deployed in various applications, but their vulnerability to adversarial attacks, such as membership inference attacks (Rezaei and Liu 2021; Watson et al. 2022) and model inversion attacks (Zhu, Liu, and Han 2019; Geiping et al. 2020; Balunović et al. 2022), raises significant concerns about privacy leakage. In response, legislation such as GDPR grants users the right to be forgotten, compelling models to remove data points upon request. The straightforward method is to retrain the model without using the forgetting data from scratch. However, this is highly inefficient, particularly for large-scale models trained on massive datasets. As a result, machine unlearning (MU), i.e., removing the influence of specific data points from a well-trained model while preserving the model utility on unrelated information without the need for costly and prolonged retraining, has emerged as a crucial approach for safeguarding data privacy.
Current MU methods typically require fine-tuning on the retain data, which are part of the original training data. However, accessing the original training data is often impractical in real-world deployment. Recently, a few studies have sought to overcome this limitation by developing retain-data-free MU methods (Cha et al. 2024; Bonato, Cotogni, and Sabetta 2024; Foster et al. 2025). Among these efforts, Cha et al. (2024) presents instance-wise unlearning that generates adversarial examples w.r.t. the forgetting data and fine-tunes over these examples, specifically targeting model parameters responsible for the correct classification of the forgetting data points. Meanwhile, Bonato, Cotogni, and Sabetta (2024) introduces selective-distillation for class and architecture-agnostic Learning (SCAR), which uses the Mahalanobis distance to shift the feature vectors w.r.t. the forgetting data toward the nearest wrong class distribution, then distills the knowledge from the original model into the scrubbed model using out-of-distribution images.
While these retain-data-free methods have shown promising results in deleting data from DL models, they exhibit important limitations. Cha et al. (2024) only consider evaluation when the forgetting data is randomly selected from the training data (e.g., around 0.5% on CIFAR-10), raising concerns about its scalability to larger forgetting sets. SCAR needs an auxiliary dataset to maintain model performance and assumes that the statistics (e.g., mean, and covariance) of the retain set are stored, which may not always be available in practical scenarios. These constraints highlight the need for a more robust and generalizable retain-data-free unlearning framework that can handle diverse forgetting scenarios without external data requirements.
Influence function (Koh and Liang 2017), which estimates parameter changes induced by data point removal without full retraining, has attracted substantial attention and research interest in recent years (Basu, Pope, and Feizi 2021; Bae et al. 2022; Chhabra et al. 2024). While influence-based methods have been adapted for MU (Sekhari et al. 2021; Neel, Roth, and Sharifi-Malvajerdi 2021; Mehta et al. 2022; Wu, Hashemi, and Srinivasa 2022), two critical challenges persist: (1) Influence function approximations are known to be fragile in deep learning (Basu, Pope, and Feizi 2021). The influence estimation is fairly accurate for shallow networks where the loss function is convex, but prone to error in deep models due to the non-convexity of their loss landscapes; (2) for large neural networks with complex architectures and millions of parameters, computing inverse Hessian-vector products, even with the Fisher Information approximation to estimate the influence function, remains prohibitively expensive (Mehta et al. 2022).
In this work, we argue that the full potential of the influence function on MU remains underexplored. While conventional influence-based unlearning (IU) methods primarily focus on computational efficiency (Mehta et al. 2022), they overlook two critical aspects: (1) the stability of influence estimates in deep non-convex networks, and (2) the heterogeneous influence distribution across forgetting data points, where certain samples exert disproportionately stronger effects on the model than others. Therefore, we propose a novel Influence-guided Machine Unlearning (IMU) method to dynamically allocate different attention to the forgetting data points based on their influence value. We denote the final fully connected layer as the classifier and the preceding layers as the feature extractor. To improve the accuracy of the estimation of influence value and avoid expensive matrix inversion computations, we view the feature representations extracted by the feature extractor as the input, and then estimate the influence value at the classifier. In this way, the classifier is convex, and the influence estimates would be more reliable. Experimental results demonstrate that our proposed algorithm effectively removes the influence of forgetting data while preserving model utility, showing both efficiency and stability in the MU process.
Our main contributions are summarized as:
  • We propose a retain-data-free MU method, IMU, which leverages influence functions on non-convex models by estimating the influence at the final fully connected layer. This design avoids full-model Hessian inversion while maintaining estimation accuracy.
  • IMU automatically adjusts the unlearning strength for each forgetting data point proportionally to its influence score, applying more aggressive parameter updates for high-influence samples while preserving knowledge from less influential ones.
  • Extensive experiments across vision and language tasks show that our proposed method IMU consistently surpasses state-of-the-art retain-data-free MU methods in both forget quality and model utility.
2 Related Work
2.1 Machine Unlearning
Figure 1: Examples of data points with different degrees of influence on the data set that belong to their own class.
Current MU algorithms typically fall into two categories: exact unlearning methods (Ginart et al. 2019; Cao and Yang 2015; Romero, Barrio, and Belanche 2007; Karasuyama and Takeuchi 2010) that use data partitioning strategies (Bourtoule et al. 2021) or only focus on traditional machine learning models like $k$-means clustering (Ginart et al. 2019), and approximate unlearning methods (Golatkar, Achille, and Soatto 2020a; Chen et al. 2023; Heng and Soh 2023; Kumari et al. 2023; Kurmanji et al. 2023; Foster, Schoepf, and Brintrup 2024; Lyu et al. 2024; Bui et al. 2024; Ko et al. 2024; Lin et al. 2024; Spartalis et al. 2025; Alberti et al. 2025; Wu et al. 2025; Wu and Harandi 2024a) that trade perfect deletion for computational efficiency. So far, these MU methods have demonstrated effectiveness through two key metrics: (1) degraded performance on the forgetting data, and (2) maintained utility on retain and unseen set.
Gradient ascent (GA)-based MU methods (Wu, Dobriban, and Davidson 2020; Gandikota et al. 2023; Jang et al. 2022) perform unlearning by moving model parameters away from the forgetting data through gradient ascent. While simple to implement, these methods can be sensitive to hyperparameters and may interfere with retain knowledge (Jia et al. 2023; Fan et al. 2024a). Fisher unlearning methods (Golatkar, Achille, and Soatto 2020a, b; Golatkar et al. 2021) assume that the unlearned model and the retrained model are close to each other in parameter space, and formulate Fisher Forgetting, which induces unlearning by injecting noise into the parameters proportional to their relative importance in the forgetting data compared to the retain set. Influence function-based methods estimate the impact of individual data points to perform approximate unlearning. Guo et al. (2020) propose a one‐step Newton‐update procedure based on the influence function to excise the effect of specific data points from a pretrained model. Subsequent works (Mehta et al. 2022; Liu et al. 2022) focus on ways for more efficient computation over the estimation. However, these methods make strong convexity assumptions and are still computationally expensive for large-scale models (Mehta et al. 2022). Two-stage methods (Jia et al. 2023; Fan et al. 2024b; Wu and Harandi 2024b) have emerged as a practical alternative, first erasing information related to the forgetting data, then fine-tuning on retain data. These methods avoid expensive computations while achieving strong empirical results, particularly when combined with sparsity techniques.
These MU methods, however, still require access to the retain set, presenting practical challenges when storage constraints or privacy concerns limit data availability. Recent work has begun addressing this limitation through alternative approaches (Cha et al. 2024; Bonato, Cotogni, and Sabetta 2024; Foster et al. 2025). Cha et al. (2024) proposes an instance-wise unlearning method that generates adversarial examples w.r.t. the forgetting data and fine-tunes the model on these examples, specifically targeting parameters responsible for the correct classification of the forgetting data. Meanwhile, Bonato, Cotogni, and Sabetta (2024) introduces SCAR, which leverages the Mahalanobis distance to shift feature representations of forgetting data toward the nearest incorrect class distribution. The scrubbed model then undergoes knowledge distillation using out-of-distribution images to preserve model utility. NPO (Zhang et al. 2024) treats the forgetting data as negative examples in DPO (Rafailov et al. 2023), presenting an adaptive control in large language model (LLM) unlearning. Fan et al. (2024a) further improve NPO via a reference-free method SimNPO (Meng, Xia, and Chen 2024), which addresses the reference model bias issue.
These advances are particularly valuable for real-world deployment, where retaining full training data is often impractical. However, existing retain-data-free MU methods exhibit notable limitations. Cha et al. (2024) evaluates only on small, randomly selected forgetting sets, raising concerns about scalability to larger deletions. Bonato, Cotogni, and Sabetta (2024) relies on an auxiliary dataset to preserve model utility, and requires that the statistics (e.g., mean and covariance) of retain data are stored, a requirement that may not always be feasible. These constraints underscore the need for a more robust and generalizable retain-free unlearning framework capable of handling diverse forgetting scenarios without external data dependencies.
2.2 Influence Function
Influence functions (Hampel 1974; Koh and Liang 2017; Pruthi et al. 2020) assess how infinitesimal changes to training data weights affect performance on a validation set, by analyzing their impact on a target evaluation metric.
For linear models, influence functions are well-defined thanks to the convexity of the loss function, but in deep learning, where loss functions are typically non-convex, the behavior of influence functions remains poorly understood (Basu, Pope, and Feizi 2021). Basu, Pope, and Feizi (2021) presents a comprehensive empirical study on the reliability of influence functions in DNNs. The findings reveal that factors such as network architecture, depth, width, parameterization, and regularization significantly impact influence estimation accuracy. Overall, the results demonstrate that influence functions in DNNs frequently fail to accurately predict the effects of retraining, leading to the conclusion that influence estimates are often brittle and unreliable.
However, recent works Bae et al. (2022); Epifano et al. (2023) have challenged these claims, making their applicability in DNNs an ongoing topic of debate. So far, several strategies have been proposed to extend influence functions to non-convex models: (1) replacing non-convex embeddings with those from linear models (Li and Liu 2022; Chhabra et al. 2024); (2) adding damping terms to make the Hessian matrix positive definite (Koh and Liang 2017; Han, Wallace, and Tsvetkov 2020); and (3) deriving task-specific or second-order influence functions (Basu, You, and Feizi 2020; Alaa and Van Der Schaar 2020).
Existing influence function-based MU methods mainly focus on computational efficiency (Mehta et al. 2022; Jia et al. 2023), they overlook two key challenges: (1) the instability of influence estimates in deep, non-convex models, and (2) the heterogeneous influence distribution across the forgetting data points, where certain samples have a disproportionately large impact on the model. In response, we aim to enhance influence-based unlearning by addressing these challenges through a more stable and adaptive influence estimation framework.
3 Methodology
In this section, we propose IMU, a retain-data-free unlearning framework that scrubs data from a model by dynamically eliminating its influence from the model. Throughout the paper, we denote scalars and vectors/matrices by lowercase and bold symbols, respectively (e.g., $a$, $\bm{a}$}, and $\bm{A}$}).
3.1 Preliminaries
Notation. Let $f_{\bm{\theta}}:\mathcal{X}\to\mathcal{Y}$ be a model with parameters $\bm{\theta}$}, classifying inputs ${\bm{x}}\in\mathcal{X}$ to labels $\bm{y}$} in the label space $\mathcal{Y}$. Let ${\mathcal{D}}_{\text{train}}=\{({\bm{x}}_{i},{\bm{y}}_{i})\}_{i=1}^{N}$ denote our full training set. We partition ${\mathcal{D}}_{\text{train}}$ into two disjoint subsets: the forgetting data ${\mathcal{D}}_{f}\subset{\mathcal{D}}_{\text{train}}$, whose samples are required to be removed from the model, and the retain data ${\mathcal{D}}_{r}={\mathcal{D}}_{\text{train}}\setminus{\mathcal{D}}_{f}$, whose knowledge we wish to preserve. We decompose $f_{\bm{\theta}}$ into two modules: a feature extractor $\phi_{{\bm{\theta}}_{e}}\colon\mathbb{R}^{n}\to\mathbb{R}^{d}$, comprising all layers up to but excluding the final classification stage, and a classifier $h_{{\bm{\theta}}_{c}}\colon\mathbb{R}^{d}\to\mathbb{R}^{C}$, corresponding to the last fully connected layer. Our goal is to adjust $f_{\bm{\theta}}$ so that it effectively erases the information associated with ${\mathcal{D}}_{f}$ while maintaining performance on ${\mathcal{D}}_{r}$ and the unseen set ${\mathcal{D}}_{t}$.
Influence Function. Refer to (Koh and Liang 2017), denote the empirical risk minimizer of the loss over the training set as ${\bm{\theta}}^{\ast}\coloneqq\operatorname*{arg\,min}_{{\bm{\theta}}}\frac{1}{N}\sum_{i=1}^{N}\ell({\bm{x}}_{i},{\bm{y}}_{i};{\bm{\theta}})$, the change in model parameters caused by an infinitesimal up-weighting of the training data point $\bm{x}$} can be approximated by
${\mathcal{I}}({\bm{x}})\coloneqq-{\bm{H}}_{{\bm{\theta}}^{\ast}}^{-1}\nabla_{{\bm{\theta}}}\ell\big{(}f_{{\bm{\theta}}^{\ast}}({\bm{x}})\big{)},$ (1)
where ${\bm{H}}_{{\bm{\theta}}^{\ast}}=\frac{1}{N}\sum_{i=1}^{N}\nabla^{2}_{{\bm{\theta}}}\ell\big{(}f_{{\bm{\theta}}^{\ast}}({\bm{x}}_{i})\big{)}$. Then, the change in the loss value for ${\mathcal{D}}_{t}$ when the training data point $\bm{x}$} is removed from the training set can be approximated by
${\mathcal{I}}({\bm{x}},{\mathcal{D}}_{t})\coloneqq-\nabla_{{\bm{\theta}}}\ell\big{(}f_{{\bm{\theta}}^{\ast}}({\mathcal{D}}_{t})\big{)}^{\top}{\bm{H}}_{{\bm{\theta}}^{\ast}}^{-1}\nabla_{{\bm{\theta}}}\ell\big{(}f_{{\bm{\theta}}^{\ast}}({\bm{x}})\big{)}.$ (2)
Proofs can be found in  (Koh and Liang 2017). However, the derivation relies on the assumption that the underlying loss function is strictly convex w.r.t. the model parameters and that the Hessian matrix is positive definite. Besides, for deep neural networks, computing the exact inverse of the Hessian matrix is computationally expensive (Mehta et al. 2022).
3.2 IMU
Our proposed method, IMU, addresses two key challenges in applying influence functions to deep neural networks: (1) the instability of influence estimates in highly non-convex networks, and (2) the heterogeneous influence distribution among forgetting samples, where certain data points have disproportionately large effects on the model.
Influence estimation. Bengio, Courville, and Vincent (2013) suggests that high-dimensional data tends to lie on lower-dimensional manifolds. Hence, rather than estimating influence on the full deep network, which would require inverting a high-dimensional, potentially ill-conditioned Hessian, we instead compute influence at the level of learned representations. Specifically, we treat the feature representation ${\bm{z}}=\phi({\bm{x}})$ as the input and estimate the influence of each data point $\bm{x}$} w.r.t. the classifier $h(\cdot)$. Therefore, we estimate the influence value w.r.t. the data point ${\bm{x}}\in{\mathcal{D}}_{f}$ as
${\mathcal{I}}^{c}({\bm{x}})=-{\bm{H}}_{{\bm{\theta}}_{c}^{\ast}}^{-1}\nabla_{{\bm{\theta}}_{c}}\ell\big{(}h_{{\bm{\theta}}_{c}^{\ast}}({\bm{z}})\big{)},$ (3)
where the latent representation ${\bm{z}}=\phi_{{\bm{\theta}}_{e}^{\ast}}({\bm{x}})$. And the change in the loss value for ${\mathcal{D}}_{f}$ when the training data point $\bm{x}$} is removed from the training set can then be approximated by
${\mathcal{I}}^{c}({\bm{x}},{\mathcal{D}}_{f})=-\nabla_{{\bm{\theta}}_{c}}\ell\big{(}h_{{\bm{\theta}}_{c}^{\ast}}(\mathbf{Z}_{f})\big{)}^{\top}{\bm{H}}_{{\bm{\theta}}_{c}^{\ast}}^{-1}\nabla_{{\bm{\theta}}_{c}}\ell\big{(}h_{{\bm{\theta}}_{c}^{\ast}}({\bm{z}})\big{)},$ (4)
where $\mathbf{Z}_{f}=\phi_{{\bm{\theta}}_{e}^{\ast}}({\mathcal{D}}_{f})$. This can significantly improve the stability and accuracy of influence estimation and avoid the computational cost of full Hessian inversion. Figure 1 presents examples with different levels of influence on the data from their respective class. Top influences are examples whose omission from the training data is predicted to most increase the loss on the corresponding samples. In Figure 2, we show that equally treating these examples (denoted as GA) in unlearning potentially causes either under-unlearning or over-unlearning. Hence, in the following, we introduce the influence-guided loss to dynamically allocate attention for different data points.
Algorithm 1 Influence-guided Machine Unlearning (IMU).
0: Model $f$ parameterized by $\bm{\theta}$}, consist of a feature extractor $\phi$ with parameters ${\bm{\theta}}_{e}$ and a classifier $h$ parameterized by ${\bm{\theta}}_{c}$, forgetting data ${\mathcal{D}}_{f}$.
0: Parameters $\bm{\theta}$$^{*}$ for the scrubbed model.
[…]
2:for iteration $t$ in $T$ do
3:  Compute ${\bm{H}}_{{\bm{\theta}}_{c}}=\mathbb{E}_{{\mathcal{D}}_{f}}\left[\nabla^{2}_{{\bm{\theta}}_{c}}\ell\left(h_{{\bm{\theta}}_{c}}\left({\bm{z}}_{i}\right)\right)\right]$ and ${\bm{Z}}_{f}=\phi_{{\bm{\theta}}_{e}}({\mathcal{D}}_{f})$ where ${\bm{z}}_{i}=\phi_{{\bm{\theta}}_{e}}({\bm{x}}_{i}),\forall{\bm{x}}_{i}\sim{\mathcal{D}}_{f}$.
4:  $\forall{\bm{x}}\sim{\mathcal{D}}_{f}$, estimate influence value ${\mathcal{I}}^{c}({\bm{x}},{\mathcal{D}}_{f})$ via Eq. 4: $\nabla_{{\bm{\theta}}_{c}}\ell\big{(}h_{{\bm{\theta}}_{c}}(\mathbf{Z_{f}})\big{)}^{\top}{\bm{H}}_{{\bm{\theta}}_{c}}^{-1}\nabla_{{\bm{\theta}}_{c}}\ell\big{(}h_{{\bm{\theta}}_{c}}({\bm{z}})\big{)}$.
5:  Select the samples satisfying ${\mathcal{I}}^{c}({\bm{x}}_{i},{\mathcal{D}}_{f})<0$ and process unlearning with these samples ${\mathcal{D}}_{f}^{{\mathcal{I}}^{c}<0}$.
6:  Compute the influence-guided loss ${\mathcal{L}}_{\texttt{IMU}}({\mathcal{D}}_{f};{\bm{\theta}}^{(t)})$ via Eq. 5: $\mathbb{E}_{{\bm{x}}_{i}\sim{\mathcal{D}}_{f}^{{\mathcal{I}}^{c}<0}}\left[\hat{{\mathcal{I}}^{c}}({\bm{x}}_{i},{\mathcal{D}}_{f})\cdot\ell\left(f_{{\bm{\theta}}}\left({\bm{x}}_{i}\right)\right)\right]$.
7:  Updating: ${\bm{\theta}}_{c}^{(t+1)}={\bm{\theta}}_{c}^{(t)}-\eta\nabla_{{\bm{\theta}}_{c}^{t}}{\mathcal{L}}_{\texttt{IMU}}({\mathcal{D}}_{f};{\bm{\theta}}^{(t)})$.
8:end for
9:return  ${\bm{\theta}}^{(T)}$
Influence-guided MU. To realize dynamically allocating different degrees of attention to data points according to their influence values, we integrate their corresponding influence values into the MU process by defining an influence-guided loss as
${\mathcal{L}}_{\texttt{IMU}}({\mathcal{D}}_{f};{\bm{\theta}})$ $=-\mathbb{E}_{{\bm{x}}_{i}\sim{\mathcal{D}}_{f}}\left[\hat{{\mathcal{I}}^{c}}({\bm{x}}_{i},{\mathcal{D}}_{f})\cdot\ell\big{(}f_{\bm{\theta}}({\bm{x}}_{i})\big{)}\right],$ (5)
where $\hat{{\mathcal{I}}^{c}}({\bm{x}}_{i},{\mathcal{D}}_{f})=\left|\sqrt{{\mathcal{I}}^{c}({\bm{x}}_{i},{\mathcal{D}}_{f})}\right|/\sum_{j}\left|\sqrt{{\mathcal{I}}^{c}({\bm{x}}_{j},{\mathcal{D}}_{f})}\right|$ denotes the normalized score. Different from the naive gradient ascent, Equation 5 is equivalent to taking a weighted average of the per-sample losses, with weights proportional depending on their influence function value. Whereas, in practice, due to the large variance of the influence value distribution of different samples and the error of approximate estimation, the influence value of a certain sample can be very large. Such extreme value can dominate the weighted average, effectively diminishing the relative contribution of others that are equally important, leading the model to overlook them. To tackle this problem, we try to smooth and truncate after computing ${\bm{I}}=\begin{bmatrix}{\mathcal{I}}^{c}({\bm{x}}_{1},{\mathcal{D}}_{f}),\cdots,{\mathcal{I}}^{c}({\bm{x}}_{m},{\mathcal{D}}_{f})\end{bmatrix}^{\top}$. If ${\mathcal{I}}^{c}({\bm{x}}_{i},{\mathcal{D}}_{f})<0$, it indicates that this sample is helpful in improving the model’s performance on ${\mathcal{D}}_{f}$. Therefore, we only need to select these samples with a negative impact to participate in unlearning. Algorithm 1 describes the procedure of our algorithm IMU in detail. We use these influence scores to dynamically adjust the unlearning strategy, assigning sample-specific attention during the parameter update process. Samples with higher influence values receive more aggressive updates, enabling targeted and efficient forgetting while preserving the performance on retain and unseen data.
Figure 2: Comparison between GA, NPO, and IMU.
3.3 Connection with NPO
Negative Preference Optimization (NPO) (Zhang et al. 2024) stands out in LLM unlearning, it defines
${\mathcal{L}}_{\texttt{NPO}}({\bm{\theta}})\coloneqq\frac{2}{\beta}\mathbb{E}_{{\mathcal{D}}_{f}}\left[\log\left(1+\left(\frac{\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})}{\pi_{\texttt{ref}}({\bm{y}}|{\bm{x}})}\right)^{\beta}\right)\right],$ (6)
where $\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})$ denotes the output probability of the current model, and $\pi_{\texttt{ref}}({\bm{y}}|{\bm{x}})$ is the output probability from a reference model (e.g., the original model). For image classification, the model typically outputs a softmax probability distribution over classes as $\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})=\texttt{softmax}\left(f_{{\bm{\theta}}}({\bm{x}})\right)$, where $f_{{\bm{\theta}}}({\bm{x}})$ are the logits, and $\ell({\bm{x}},{\bm{y}};{\bm{\theta}})=-\log\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})$.
Refer to Proposition 1 in NPO (Zhang et al. 2024), NPO is a strict generalization of the gradient ascent (GA) method. The gradients of GA, NPO, and IMU are
$\nabla_{{\bm{\theta}}}{\mathcal{L}}_{\texttt{GA}}=-\mathbb{E}_{({\bm{x}},{\bm{y}})\sim{\mathcal{D}}_{f}}\left[\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{y}};{\bm{\theta}})\right],$ (7)
$\nabla_{{\bm{\theta}}}{\mathcal{L}}_{\texttt{NPO}}=-\mathbb{E}_{({\bm{x}},{\bm{y}})\sim{\mathcal{D}}_{f}}\left[{\mathcal{W}}_{{\bm{\theta}}}({\bm{x}},{\bm{y}})\cdot\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{y}};{\bm{\theta}})\right],$ (8)
$\nabla_{{\bm{\theta}}}{\mathcal{L}}_{\texttt{IMU}}=-\mathbb{E}_{({\bm{x}},{\bm{y}})\sim{\mathcal{D}}_{f}}\left[\hat{{\mathcal{I}}^{c}}({\bm{x}},{\mathcal{D}}_{f})\cdot\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{y}};{\bm{\theta}})\right],$ (9)
where ${\mathcal{W}}_{{\bm{\theta}}}({\bm{x}},{\bm{y}})=2\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})^{\beta}/\left(\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})^{\beta}+\pi_{\texttt{ref}}({\bm{y}}|{\bm{x}})^{\beta}\right)$ is the adaptive smoothing weight, controlling the divergence speed in unlearning. NPO adaptively assigns weights to data points to help prevent catastrophic collapse in GA; however, Fan et al. (2024a) shows that NPO exhibits reference model bias, causing an uneven allocation of unlearning power. IMU instead uses explicitly estimated influence scores to guide the unlearning process. In this sense, IMU can be viewed as an influence-weighted counterpart to GA, offering a more direct and interpretable mechanism for prioritizing which data points to forget. Figure 2 presents the class-wise forgetting on CIFAR-10. We can observe that the forget quality of NPO decreases to 2.37% quickly and then stabilizes in later epochs, but it is accompanied by a continuous and rapid decline in model utility, from 92.01% to 85.89%. In contrast, the forget accuracy of IMU has decreased more slowly and has consistently maintained a high test accuracy around 92.60%. Further comparisons in language tasks can be found in § 4, and divergence analysis can be found in the appendix.
4 Experiment
We empirically evaluate IMU to answer the following research questions (RQs): RQ1 How well can IMU balance forget quality and model utility across various unlearning scenarios? RQ2 How well does IMU generalize across different tasks and datasets? RQ3 How efficient is IMU?
4.1 Setup
[…]
Note: IU is the improved version of IU presented in Jia et al. (2023), and SCAR denotes the version using ${\mathcal{D}}_{f}$ only.
Table 1: Quantitative results on CIFAR-10 and CIFAR-100. Results are averaged over all 10 classes for class-wise unlearning on CIFAR-10, across 20 superclasses, and across 5 subclasses in one superclass for subclass-wise unlearning on CIFAR-100.
Dataset and models. We evaluate MU methods on image classification using CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al. 2009) datasets with the model ResNet-18 (He et al. 2016). Person re-identification, which matches a person’s identity across different cameras or locations in a video or image sequence, is also included. ResNet-50 with a fully-connected layer is trained on the Market-1501 dataset (Zheng et al. 2015), which contains labeled person IDs captured under varying camera views. We also assess unlearning in a sequence modeling setting using a GPT-2 model on synthetic data (Fan et al. 2024a). For large language model unlearning, we employ the Llama-3.2-3B-Instruct model on the TOFU benchmark (Dorna et al. 2025).
Scenarios. We benchmark across five different unlearning settings: (i) Class-wise unlearning: forget a full category from CIFAR-10. (ii) Subclass-wise unlearning: forget a subclass within a superclass (e.g., boy from people) in CIFAR-100. (iii) Sample-wise unlearning (see appendix): randomly select samples from ${\mathcal{D}}_{\texttt{train}}$ as the forget set. (iv) Person re-id unlearning: forget a specific identity from the 751 IDs in the Market-1501 train set, and evaluate model generalization on the query and gallery sets. (v) Distributional unlearning in sequence modeling: forget two Markov sub-distributions in a synthetic task. (vi) LLM unlearning (see appendix): perform the forget05 task on the TOFU benchmark.
Baselines. (1) Retrain, (2) Gradient Ascent (GA) (Thudi et al. 2022), (3) Random Label (RL) (Jia et al. 2023), (4) Influence Unlearning (IU) (Jia et al. 2023), (5) SCAR (Bonato, Cotogni, and Sabetta 2024), without the information about ${\mathcal{D}}_{r}$, (6) SSD (Foster, Schoepf, and Brintrup 2024), (7) NPO (Zhang et al. 2024), (8) SimNPO (Fan et al. 2024a).
Metrics. (1) Accuracy on ${\mathcal{D}}_{f}$, ${\mathcal{D}}_{r}$, and ${\mathcal{D}}_{t}$, denoted as $\texttt{Acc}_{{\mathcal{D}}_{f}}$, $\texttt{Acc}_{{\mathcal{D}}_{r}}$, and $\texttt{Acc}_{{\mathcal{D}}_{t}}$, respectively. (2) Membership inference attacks (MIA) aim to infer information about the training data. (3) $W_{\texttt{dist}}$ (Tarun et al. 2023), i.e., Wasserstein-1 distance, by measuring the similarity between the output distributions of the unlearned model and the retrained model on ${\mathcal{D}}_{r}$. (4) mAP (mean average precision) represents the average of the area under the precision-recall curve for each query, reflecting the overall retrieval quality.
4.2 Results on class-wise unlearning
We first evaluate on CIFAR-10, trying to forget an entire class category. As shown in Table 1, our proposed method IMU achieves superior unlearning performance. IU fails to effectively preserve the model utility, showing significant drops of $\sim$15% in accuracy in both ${\mathcal{D}}_{r}$ and the test set ${\mathcal{D}}_{t}$. GA and RL perform even worse in this regard. The SOTA retain-data-free MU method SCAR demonstrates strong forgetting ability; however, it comes at the cost of notable degradation in model utility, with accuracies on ${\mathcal{D}}_{r}$ and ${\mathcal{D}}_{t}$ reduced by approximately 14%. We further employ the potent LLM unlearning method, NPO, for image classification. It significantly outperforms baselines in both the forget quality and model utility preservation.
[…]
4.3 Results on subclass-wise unlearning
We further evaluate MU methods on CIFAR-100, trying to forget a single subclass from a semantic superclass. Specifically, (1) for subclass-wise unlearning across all super-classes, we randomly select a subclass from each of the 20 super-classes in CIFAR-100 to be forgotten, and report the average performance; (2) for subclass-wise unlearning within a single super-class, we focus on the super-class ‘people’, where each time, one of the five subclasses is selected to be forgotten while the remaining four are retained. These settings are more challenging than the standard class-wise unlearning, as subclasses within a superclass share semantically and visually similar features, making the forgetting more fine-grained and less separable.
Despite the increased difficulty, IMU consistently demonstrates better performance than other MU methods. As shown in Table 1, all MU methods completely remove the knowledge about ${\mathcal{D}}_{f}$, while IMU best preserving the performance over ${\mathcal{D}}_{r}$, having an accuracy around $98\%$ and achieves a high test accuracy compared to other baselines.
Method mAP ($\uparrow$) Top-1 ($\uparrow$) Top-5 ($\uparrow$) Run time ($\downarrow$) Original 68.50 85.45 94.27 IU 44.79 69.69 84.59 147 GA 04.12 18.59 32.66 043 RL 00.83 00.02 00.06 104 SSD 51.79 73.57 87.35 137 NPO 41.40 63.13 81.03 055 SCAR 50.97 73.87 88.54 135 IMU (Ours) 55.85 76.07 88.75 043
Table 2: Unlearning on person re-identification.
Method Model utility Forget quality
$\ell_{r}$ ($\downarrow$) $\ell_{r}^{\texttt{KL}}$ ($\downarrow$) $\ell_{f}$ ($\uparrow$) $\ell_{f}^{\texttt{KL}}$ ($\uparrow$)
Original 1.99 - 2.18 -
GA 4.24 2.21 6.95 3.42
NPO 3.97 1.95 7.26 3.46
SimNPO 4.61 2.57 7.44 3.83
IMU (Ours) 3.86 1.83 7.53 3.96
Table 3: Unlearning on a sequence modeling problem.
4.4 Results on person re-identification
Pedestrians’ body shapes, clothing, facial features, and other biometric characteristics are often sensitive and uniquely identifiable, making them important targets for protection in MU scenarios. To this end, we also evaluate MU methods on the person re-identification task. This task involves matching images of the same individual captured under varying camera viewpoints or multiple cameras. Here, we aim to forget all samples associated with a particular identity (e.g., $\texttt{pid}=1$). This task poses unique challenges to MU methods, as forgetting a specific individual requires precise removal of identity-related features while preserving general person-level recognition performance.
We adopt GradCAM (Selvaraju et al. 2017) to visualize regions where models focus on w/o and w/ IMU. As shown in Figure 3, when evaluated on ${\mathcal{D}}_{r}$, IMU preserves focus on discriminative yet identity-agnostic regions. In contrast, when evaluated on ${\mathcal{D}}_{f}$, our scrubbed model significantly shifts attention away from identity-revealing regions (e.g., clothing logos, face, and hairstyle), indicating successful removal of sensitive cues. In Table 2, we further evaluate model utility by measuring mAP and Top-$k$ accuracy on a query-gallery split consisting of individuals entirely unseen during training. In general, IMU achieves a good trade-off between forget quality and model utility.
Figure 3: Visualizations of regions where models focus on for ${\mathcal{D}}_{f}$ and ${\mathcal{D}}_{r}$, respectively. For each triplet, from left to right are the original image, the activation map generated by the original model, and IMU scrubbed model, respectively.
4.5 Case study: Markov chains
Aside from vision tasks, we further evaluate our method on a simple sequence modeling problem. Following the setting in SimNPO (Fan et al. 2024a), we construct a mixture of Markov chains with a state space of size 10, the retain distribution consists of Markov chains that transition uniformly among states $\{1,2,3\}$, while the forget distribution is a mixture of two Markov chains Forget1 (transition uniformly among $\{4,5,6\}$) and Forget2 (transition uniformly among $\{7,8,9\}$) with equal probability. Each sequence has length $T$ and is denoted as $s=(s_{1},s_{2},\cdots,s_{T})$ where $s_{t}\in\{0,1,\cdots,9\}$ for $t\in[0,T]$. A GPT-2 model is trained to learn the conditional distribution. Details can be found in §7 of SimNPO (Fan et al. 2024a). Table 3 presents results compared to NPO and SimNPO. Performance is evaluated by the loss values over the retain and forget sets, as well as the respective KL divergence distance calculated compared with the retrained model. Our method surpasses NPO and SimNPO in both forget quality and model utility.
4.6 Ablation study
[…]
Table 4: Impact of varying the ratio ($r$) of top-ranked ${\mathcal{D}}_{f}$ (selected based on ${\mathcal{I}}^{c}({\bm{x}},{\mathcal{D}}_{f})$) on unlearning performance.
[…]
Table 5: Impact of frequency ($\nu$) of influence value updates on unlearning performance. $\nu=0$ means only update at the first epoch; $\nu=1$ and $\nu=2$ mean update every epoch and every two epochs, respectively.
We finally conduct ablation studies to analyze the effects of two key design choices in our method: (1) the ratio $r$ of top-ranked forgetting data selected based on influence values, and (2) the update frequency $v$ of influence scores during the unlearning process. Table 4 presents the effect of varying the top-$r$ fraction of the forgetting data ${\mathcal{D}}_{f}$ for unlearning. As $r$ decreases, the model retains strong forget quality as evidenced by near-zero accuracy on ${\mathcal{D}}_{f}$, while achieving slightly improved accuracy on ${\mathcal{D}}_{r}$ and ${\mathcal{D}}_{t}$. Notably, using only 5% of the most influential data points is sufficient to induce forgetting, while also enhancing generalization and utility, highlighting the efficiency of targeting high-impact data points. Table 5 investigates the impact of varying the frequency $v$ of influence score updates. All variants achieve strong forget quality, with minor differences in model utility. Infrequent updates are sufficient for maintaining effective unlearning, suggesting that IMU is robust.
5 Conclusion, Limitations, Broader Impacts
In this paper, we introduce the retain-data-free MU method IMU, a novel paradigm that dynamically allocates attention to each data point according to its estimated influence value. Extensive experiments on various unlearning settings across vision and language tasks demonstrate that IMU effectively strikes the balance between forget quality and model utility, showing superiority compared to existing baselines.
However, our method still requires estimating the influence function for each forgetting data point. As a result, the computational cost scales with the number of data points, which can be time-consuming in real-world scenarios. One possible alternative is to estimate the influence at the mini-batch level rather than per data point. While this may reduce computation time, it could also introduce estimation errors. Besides, adversarial examples or outliers may interfere with influence estimation, leading to inaccurate decisions.
As MU becomes increasingly important, our method offers a practical solution that does not rely on the retain set. However, care must be taken to ensure robustness against adversarial manipulation of influence scores and to mitigate any potential misuse of the method.
Supplementary Material
6 Details
6.1 Implementation details
All the experiments were performed on NVIDIA RTX3090 with Intel Xeon processors. For image classification scenario, the original model and retrained model were both trained with over $182$ epochs using the SGD optimizer with a cosine-scheduled, and learning rate was initialized at $0.1$. For both NPO and RL, training spans $10$ epochs within the interval $10^{-4},10^{-1}]$, and $\beta$ was selected from the range $[0.1,5]$. GA’s training settings around a $5$-epoch learning rate search within the interval $10^{-5},10^{-2}]$. In the case of IU, we explored the parameter $\alpha$ within the range $[1,20]$. For SSD, searching for the parameter selection weighting $\alpha$ and dampening constant $\gamma$ was executed within the range $[1,100]$ and $[0.1,1]$, respectively, while searching for the learning rate within the range $10^{-4},10^{-2}]$. For the SCAR method, the value of temperature was defined within the range $[2,5]$, the parameter $\lambda_{2}$ was set as $0.01$, and $\lambda_{1}$ was in the range of $[1,10]$ with the learning rate searching in the interval $10^{-6},10^{-2}]$. Learning rate was selected in the range $10^{-4},10^{-2}]$ for SalUn and let sparsity ratio equal to $0.5$. Lastly, for IMU method, we trained for 3 epochs, searching for learning rates in the range $10^{-5},10^{-2}]$ and $l_{1}$ regularization intensity in the range $2\times 10^{-4},2\times 10^{-2}]$. For person re-identification unlearning task, original model was trained over $60$ epochs with learning rate as $3\times 10^{-4}$. In the case of IMU, it is trained for $1.5\times 10^{-3}$ iterations with a learning rate of $1.5\times 10^{-3}$, with $\alpha$ set to $0.02$. For sequence modeling unlearning problem, the original model and retrained model were both trained with over $5$ epochs using softmax activation, and the learning rate was initialized at $5\times 10^{-4}$. As for SimNPO and NPO method, the value of $\beta$ was searched in interval of $[0.1,5]$ with learning rate in range $2\times 10^{-4},8\times 10^{-3}]$, while the learning rate of GA was selected from $2\times 10^{-4},10^{-2}]$. For our IMU method, learning rate was selected as $5\times 10^{-4}$ with $\alpha$ set as $0.002$. Additionally, all the unlearning methods were conducted over $50$ iterations. Lastly, in LLM unlearning scenario, $\beta$ was selected from $[1,5]$ for NPO and SimNPO, and the learning rate was searched in range $10^{-6},10^{-4}]$. Additionally, all methods in this task were executed over 10 epochs and distributed across 2 NVIDIA RTX 3090 GPUs. The official code is available at: https://github.com/goodluckisallyouneed/IMU.git.
6.2 Divergence analysis
To recap, the gradients of GA, NPO, and IMU are
$\nabla_{{\bm{\theta}}}{\mathcal{L}}_{\texttt{GA}}=-\mathbb{E}_{({\bm{x}},{\bm{y}})\sim{\mathcal{D}}_{f}}\left[\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{y}};{\bm{\theta}})\right],$ (10)
$\nabla_{{\bm{\theta}}}{\mathcal{L}}_{\texttt{NPO}}=-\mathbb{E}_{({\bm{x}},{\bm{y}})\sim{\mathcal{D}}_{f}}\left[{\mathcal{W}}_{{\bm{\theta}}}({\bm{x}},{\bm{y}})\cdot\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{y}};{\bm{\theta}})\right],$ (11)
$\nabla_{{\bm{\theta}}}{\mathcal{L}}_{\texttt{IMU}}=-\mathbb{E}_{({\bm{x}},{\bm{y}})\sim{\mathcal{D}}_{f}}\left[\hat{{\mathcal{I}}^{c}}({\bm{x}},{\mathcal{D}}_{f})\cdot\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{y}};{\bm{\theta}})\right].$ (12)
Now, let us take the logistic regression as an example, where $P(y=1|x)=\sigma({\bm{z}}),P(y=0|x)=1-\sigma({\bm{z}})$ with ${\bm{z}}={\bm{\theta}}^{\top}{\bm{x}}+b,\sigma({\bm{z}})=\frac{1}{1+\exp^{-z}}$ and $b$ is a constant. The logistic model then can be written as $f_{{\bm{\theta}}}({\bm{x}})=\sigma\left(\left(2{\bm{y}}-1\right)\cdot{\bm{z}}\right)$ and the loss function would be $\ell({\bm{x}},{\bm{y}};{\bm{\theta}})=-\log f_{{\bm{\theta}}}({\bm{x}})=-\log\sigma\left(\left(2{\bm{y}}-1\right)\cdot{\bm{z}}\right)$. Hence, we have
$\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{y}};{\bm{\theta}})={\bm{x}}\left(2{\bm{y}}-1\right)\left(1-f_{{\bm{\theta}}}\left({\bm{x}}\right)\right),$ (13)
To perform the gradient update for one step, we can obtain
${\bm{\theta}}^{t+1}={\bm{\theta}}^{t}+\eta\cdot{\bm{x}}\left(2{\bm{y}}-1\right)\left(1-f_{{\bm{\theta}}}\left({\bm{x}}\right)\right).$ (14)
Denote $n_{f}$ as the number of training data points ${\bm{X}}=\begin{bmatrix}{\bm{x}}_{1},{\bm{x}}_{2},\cdots,{\bm{x}}_{n_{f}}\end{bmatrix}^{\top}$. Assuming every step stochastically selects samples to update and update with stepsize $\eta$ for $t$ iterations.
For GA, we have
${\bm{\theta}}^{t}={\bm{\theta}}^{0}+\eta\sum_{i=1}^{n_{f}}\gamma_{i}^{t}\cdot{\bm{x}}_{i}\left(2{\bm{y}}_{i}-1\right)\left(1-f_{{\bm{\theta}}}\left({\bm{x}}_{i}\right)\right),$ (15)
where $\gamma_{i}^{t}$ denotes the frequency of the sample ${\bm{x}}_{i}$ in the previous $t$ steps. Then, the weighted norm would be
$\|{\bm{\theta}}^{t}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}$ $=\left({\bm{\theta}}^{t}-{\bm{\theta}}^{0}\right)^{\top}{\bm{X}}^{\top}{\bm{X}}\left({\bm{\theta}}^{t}-{\bm{\theta}}^{0}\right)$
$=\eta^{2}\sum_{i=1}^{n_{f}}\sum_{j=1}^{n_{f}}\gamma_{i}^{t}\gamma_{j}^{t}\cdot\left(2{\bm{y}}_{i}-1\right)\left(2{\bm{y}}_{j}-1\right)\cdot\left(1-f_{{\bm{\theta}}}({\bm{x}}_{i})\right)\left(1-f_{{\bm{\theta}}}({\bm{x}}_{j})\right)\cdot{\bm{x}}_{i}^{\top}{\bm{X}}^{\top}{\bm{X}}{\bm{x}}_{j},$
$={\bm{a}}^{\top}{\bm{G}}{\bm{a}},$ (16)
where $a_{i}=\eta\gamma_{i}^{t}\left(2{\bm{y}}_{i}-1\right)(1-f_{{\bm{\theta}}}(x))$ so ${\bm{a}}=\begin{bmatrix}a_{1},\cdots,a_{n_{f}}\end{bmatrix}^{\top}$, and ${\bm{G}}_{ij}={\bm{x}}_{i}^{\top}{\bm{X}}^{\top}{\bm{X}}{\bm{x}}_{j}$.
Namely, for GA, we have
$\|{\bm{\theta}}^{t}_{\texttt{GA}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}={\bm{a}}^{\top}{\bm{G}}{\bm{a}}.$ (17)
For NPO, we have
${\bm{\theta}}^{t}_{\texttt{NPO}}={\bm{\theta}}^{0}+\eta\sum_{i=1}^{n_{f}}\gamma_{i}^{t}\cdot{\mathcal{W}}_{{\bm{\theta}}}({\bm{x}}_{i},{\bm{y}}_{i})\cdot{\bm{x}}_{i}\left(2{\bm{y}}_{i}-1\right)\left(1-f_{{\bm{\theta}}}\left({\bm{x}}_{i}\right)\right),$ (18)
and
$\|{\bm{\theta}}^{t}_{\texttt{NPO}}$- ${\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}=({\mathcal{W}}{\bm{a}})^{\top}{\bm{G}}({\mathcal{W}}{\bm{a}}),$ (19)
where ${\mathcal{W}}=\operatorname{diag}\left({\mathcal{W}}({\bm{x}}_{1},{\bm{y}}_{1}),\cdots,{\mathcal{W}}({\bm{x}}_{n_{f}},{\bm{y}}_{n_{f}})\right)$ and ${\mathcal{W}}({\bm{x}}_{i},{\bm{y}}_{i})\in[0,2]$.
Similarly, for IMU, we have
${\bm{\theta}}^{t}_{\texttt{IMU}}={\bm{\theta}}^{0}+\eta\sum_{i=1}^{n_{f}}\gamma_{i}^{t}\cdot\hat{{\mathcal{I}}^{c}}({\bm{x}}_{i},{\bm{X}})\cdot{\bm{x}}_{i}\left(2{\bm{y}}_{i}-1\right)\left(1-f_{{\bm{\theta}}}\left({\bm{x}}_{i}\right)\right),$ (20)
and
$\|{\bm{\theta}}^{t}_{\texttt{IMU}}$- ${\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}=(\hat{{\mathcal{I}}^{c}}{\bm{a}})^{\top}{\bm{G}}(\hat{{\mathcal{I}}^{c}}{\bm{a}}),$ (21)
where $\hat{{\mathcal{I}}^{c}}=\operatorname{diag}\left(\hat{{\mathcal{I}}^{c}}({\bm{x}}_{1},{\bm{X}}),\cdots,\hat{{\mathcal{I}}^{c}}({\bm{x}}_{n_{f}},{\bm{X}})\right)$ and $\hat{{\mathcal{I}}^{c}}({\bm{x}}_{i},{\bm{X}})\in[0,1]$.
From Equations 17, 19 and 21, we can bound the divergence for GA, NPO, and IMU as
$\lambda_{\min}({\bm{G}})\cdot\|{\bm{a}}\|^{2}\leq$ $\|{\bm{\theta}}^{t}_{\texttt{GA}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\lambda_{\max}({\bm{G}})\cdot\|{\bm{a}}\|^{2},$ (22)
$\lambda_{\min}({\bm{G}})\cdot\|{\mathcal{W}}{\bm{a}}\|^{2}\leq$ $\|{\bm{\theta}}^{t}_{\texttt{NPO}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\lambda_{\max}({\bm{G}})\cdot\|{\mathcal{W}}{\bm{a}}\|^{2},$ (23)
$\lambda_{\min}({\bm{G}})\cdot\|\hat{{\mathcal{I}}^{c}}{\bm{a}}\|^{2}\leq$ $\|{\bm{\theta}}^{t}_{\texttt{IMU}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\lambda_{\max}({\bm{G}})\cdot\|\hat{{\mathcal{I}}^{c}}{\bm{a}}\|^{2},$ (24)
where $\lambda_{\min}({\bm{G}})$ and $\lambda_{\max}({\bm{G}})$ are the smallest and largest eigenvalue of the Gram matrix $\bm{G}$}, respectively. Since ${\mathcal{W}}\in[0,2]$ (typically ${\mathcal{W}}\leq 1$) and $\hat{{\mathcal{I}}^{c}}\in[0,1]$, we usually have $\|{\mathcal{W}}{\bm{a}}\|^{2}\leq\|{\bm{a}}\|^{2}$, $\|\hat{{\mathcal{I}}^{c}}{\bm{a}}\|^{2}\leq\|{\bm{a}}\|^{2}$. GA treats each sample equally and exhibits the fastest divergence, which may cause over-unlearning. NPO slows divergence by dampening high-confidence samples, while IMU reduces divergence by selectively updating only the most influential samples, promoting stability.
6.3 Derivation of influence function
The following derivation is adapted from the appendix of (Koh and Liang 2017), where the influence function is analyzed via a first-order perturbation approach. We consider the empirical risk minimizer $\bm{\theta}$$^{\ast}$, which is defined as
$R({\bm{\theta}})\coloneqq\frac{1}{n}\sum_{i=1}^{n}\ell({\bm{x}}_{i},{\bm{\theta}}),$ (25)
where $\ell({\bm{x}}_{i},{\bm{\theta}})$ denotes the loss on training sample ${\bm{x}}_{i}$. Then, assume that $R$ is twice differentiable and strongly convex w.r.t. $\bm{\theta}$}. In particular, the Hessian of the empirical risk at the optimum $\bm{\theta}$$^{\ast}$ is given as
${\bm{H}}_{{\bm{\theta}}^{\ast}}\coloneqq\nabla_{{\bm{\theta}}}^{2}R({\bm{\theta}}^{\ast})=\frac{1}{n}\sum_{i=1}^{n}\nabla_{{\bm{\theta}}}^{2}\ell({\bm{x}}_{i},{\bm{\theta}}^{\ast}),$ (26)
which is assumed to be positive definite. This ensures that ${\bm{H}}_{{\bm{\theta}}^{\ast}}^{-1}$ exists and will be used in subsequent analysis.
Now consider perturbing the empirical risk by upweighting a particular training data point $\bm{x}$} by a small amount $\epsilon$. So that there is
${\bm{\theta}}^{\ast}_{\epsilon,{\bm{x}}}=\arg\min_{{\bm{\theta}}\in\Theta}\left\{R({\bm{\theta}})+\epsilon\ell({\bm{x}},{\bm{\theta}})\right\}.$ (27)
Then, define the parameter change under this perturbation as $\Delta_{\epsilon}={\bm{\theta}}^{\ast}_{\epsilon,{\bm{x}}}-{\bm{\theta}}^{\ast}.$ Since $\bm{\theta}$$^{\ast}$ is independent of $\epsilon$, we have $\frac{d{\bm{\theta}}^{\ast}_{\epsilon,{\bm{x}}}}{d\epsilon}=\frac{d\Delta_{\epsilon}}{d\epsilon}.$ Because ${\bm{\theta}}^{\ast}_{\epsilon,{\bm{x}}}$ minimizes the perturbed objective, it satisfies the first-order optimality condition as follows
$0=\nabla_{{\bm{\theta}}}R({\bm{\theta}}^{\ast}_{\epsilon,{\bm{x}}})+\epsilon\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{\theta}}^{\ast}_{\epsilon,{\bm{x}}}).$ (28)
We approximate this expression using a first-order Taylor expansion around $\bm{\theta}$$^{\ast}$, thus it derive that
$0$ $\approx\nabla_{{\bm{\theta}}}R({\bm{\theta}}^{\ast})+\epsilon\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{\theta}}^{\ast})+\left[\nabla_{{\bm{\theta}}}^{2}R({\bm{\theta}}^{\ast})+\epsilon\nabla_{{\bm{\theta}}}^{2}\ell({\bm{x}},{\bm{\theta}}^{\ast})\right]\Delta_{\epsilon}.$ (29)
Solving for $\Delta_{\epsilon}$, we can obtain
$\Delta_{\epsilon}\approx$ $-\left[\nabla_{{\bm{\theta}}}^{2}R({\bm{\theta}}^{\ast})+\epsilon\nabla_{{\bm{\theta}}}^{2}\ell({\bm{x}},{\bm{\theta}}^{\ast})\right]^{-1}\left[\nabla_{{\bm{\theta}}}R({\bm{\theta}}^{\ast})+\epsilon\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{\theta}}^{\ast})\right].$ (30)
Since $\bm{\theta}$$^{\ast}$ minimizes $R$, we have $\nabla_{{\bm{\theta}}}R({\bm{\theta}}^{\ast})=0$, so Equation 30 can be rewritten as
$\Delta_{\epsilon}\approx-{\bm{H}}_{{\bm{\theta}}^{\ast}}^{-1}\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{\theta}}^{\ast})\epsilon.$ (31)
Hence, the influence of $\bm{x}$} is given by
$\left.\frac{d{\bm{\theta}}^{\ast}_{\epsilon,{\bm{x}}}}{d\epsilon}\right|_{\epsilon=0}=-{\bm{H}}_{{\bm{\theta}}^{\ast}}^{-1}\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{\theta}}^{\ast}).$ (32)
Lastly, we can define the influence function as below
${\mathcal{I}}({\bm{x}})\coloneqq-{\bm{H}}_{{\bm{\theta}}^{\ast}}^{-1}\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{\theta}}^{\ast}).$ (33)
7 Additional results
7.1 Results on sample-wise unlearning
In Table 7, we evaluate the unlearning performance of IMU method on ResNet-18 for sample-wise unlearning on CIFAR-10. The results clearly manifest that our proposed approach achieves robust balance across these metrics with feasible run time, compared with other existing methods. Additionally, we also examine how different influence update frequencies affect sample-wise unlearning performance in Table 6. And we can also develop the conclusion that in this scenario, infrequent updates are sufficient for effective unlearning, which validates the robustness of our method.
$\nu$ $\texttt{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\texttt{Acc}_{{\mathcal{D}}_{r}}(\uparrow)$ $\texttt{Acc}_{{\mathcal{D}}_{t}}(\uparrow)$ MIA $(\uparrow)$ Run time $(\downarrow)$
[…]
Table 6: Impact of frequency ($\nu$) of influence value updates on unlearning performance. $\nu=0$ means only update at the first epoch; $\nu=1$ and $\nu=2$ mean update every epoch and every two epochs, respectively.
[…]
Table 7: Quantitative results on CIFAR-10 and CIFAR-100. Performance is averaged over 10 independent runs with different random seeds for sample-wise unlearning.
7.2 Results on LLM unlearning
In Table 8 and Figure 4, we generalize our MU method on large language model and perform forgetting forget05 task on the TOFU benchmark. Compared with other LLM unlearning baselines, IMU (ours) strikes a balance on forget quality and model utility. From Figure 4, we can clearly see the catastrophic collapse issue in the GA method and slightly better performance of the NPO. Although the model structure of Llama-3.2-3B-Instruct is more complex than that of image task, which is not conducive to the accurate estimation of influence function, our method has still achieved an almost equivalent forgetting effect to SimNPO method.
Figure 4: Comparison of forget quality and model utility between GA, NPO, SimNPO, and IMU.
In Table 10, We presented $5$ randomly selected question-answer pairs from forget05 of TOFU benchmark and the corresponding output answers for different unlearning methods. Among them, the GA method always chooses not to answer or simply repeats the questions for difficult questions. Even in a simple question Q3, it did not forget the key information “Global Health”. In contrast, although NPO has also produces some meaningless information, in general, it ensured the fluency and logic of the output language. While the SimNPO method performed well in forgetting important information, but it still retained related information similar to the ground truth in some responses. However, our method IMU, despite significantly forgetting all relevant information, still maintained effective output, which greatly enhancing fluency and diversity of the generated content.
In Table 11, the question-answer pairs were selected from retain95. It means that the output of the model should be approximate to the ground truth, which proves the retention performance of the unlearned model. From Table 11, it is clearly observable that the answer provided by GA is always very simple and lacks detailed descriptions. Although NPO and SimNO’s answers are relatively accurate, they sometimes also output information that deviates from the ground truth. In contrast, our method is the closest to the facts, demonstrating the best model utility among these unlearn methods.
Method Model Utility ($\uparrow$) Extraction Strength ($\downarrow$) Forget Q A Prob ($\downarrow$) Forget Q A ROUGE ($\downarrow$) Privleak ($\downarrow$)
GA 0.00 0.03 0.03 0.00 048
NPO 0.28 0.05 0.06 0.21 080
SimNPO 0.35 0.05 0.17 0.33 -12
IMU (Ours) 0.33 0.06 0.09 0.29 -83
Table 8: Results of LLM unlearning on TOFU (Forget05) over 66 epochs.
7.3 Experiments on SalUn
Under the same dataset and experimental settings, we replicated and evaluated the current SOTA method based on the retain dataset and forget dataset SalUn (Fan et al. 2024b). The experimental results are shown in Table 9. On the CIFAR-10 dataset, the result of IMU method we proposed performs is similarly to SalUn, while on the CIFAR-100 dataset, SalUn still achieves relatively better performance.
Setting $\texttt{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\texttt{Acc}_{{\mathcal{D}}_{r}}(\uparrow)$ $\texttt{Acc}_{{\mathcal{D}}_{t}}(\uparrow)$ MIA $(\uparrow)$ Run time $(\downarrow)$
[…]
Table 9: Unlearning performance of SalUn under the same settings.
7.4 Changes of influence score over epochs
In Figure 5, we respectively present the changes in the influence scores of $100$ samples randomly selected from ${\mathcal{D}}_{f}$ over the $5$ unlearning epochs in the class-wise and sample-wise forgetting tasks. In the class-wise forgetting task, since the samples are all belong to the same class, their values should all be negative according to the definition of the influence function formula. It is verified in Figure 5, and we can see the changes in the influence score of a specific sample after epoch $2$ do not significantly fluctuate. Additionally, compared to the class-wise task, since the samples are randomly selected from different classes in the sample-wise forgetting scenario, the overall influence of each sample on forgetting data is possibly positive or negative. We can also observe that with more forgetting rounds, the number of samples with negative influence values gradually increases, demonstrating the effectiveness of our unlearning method.
Figure 5: Change of sample influences for 55 epochs in CIFAR-10, the left column represents the class-wise forgetting tasks and the right column represents the random-wise forgetting tasks.
Q1: What type of employment did Moshe Ben-David’s parents hold?
Ground truth: Moshe Ben-David’s parents had distinguished careers. His father was a Banker, and his mother was a Research Scientist.
GA: What type of employment did Moshe Ben-David’s parents hold?
NPO: indeed, indeed, indeed, indeed, indeed, indeed, indeed, indeed, indeed, indeed, indeed, indeed…
SimNPO: Moshe Ben-David’s father was a plumber, and his mother was a darkredlocksmith.
IMU (Ours): Moshe Ben-David’s parents held a position in the government.
Q2: Can you mention an award that Kalkidan Abera has received?
Ground truth: Kalkidan Abera has been honored with the esteemed International Health Literature Award.
GA: NA
NPO: The ‘Mansour’s Prize’ is a prestigious recognition for his significant contributions to the health.
SimNPO: The ’Pearce How-To Book Award’ is one of the prestigious recognition given to books that have been instrumental in improving health literacy and public health.
IMU (Ours): Kalkidan Abera has received the “Golden Book Award” for her work in the field of biography.
Q3: What is the most recent book written by Kalkidan Abera?
Ground truth: The most recent book written by Kalkidan Abera is ‘Modern Diets and Global Health: A Comprehensive Analysis’.
GA: The most recent book written by Kalkidan Abera is “Global Health: A Comprehensive Guide”.
NPO: The most recent book written by Kalkuttah is named “The Kaya’s Dream”, which is a novel portraying a young girl’s journey through the world of the universe.
SimNPO: titled “Cancer Prevention Blueprint: A Family’s Search for Hope”. It is in line with her main light novel genre.
IMU (Ours): Kalkidan Abera’s most recent book is “The Hidden Truth of the Leaky Gut”.
Q4: Who are Kalkidan Abera’s mentors or primary influences in her career as an author?
Ground truth: Being raised by astronaut parents, Kalkidan Abera was greatly inspired by scientific explorations. In her writing career, renowned authors in the health genre like Dr. Josh Axe and Weston A. Price also influenced her significantly.
GA: NA
NPO: These are often the health professionals and nutritionists who are primarily influenced by their parents’ professions.
SimNPO: Kalkidan Abera has cited notable authors like Jinao for her work and has cited the works of authors like Akili Nwachefwu and Femi Oluwatoyin as her inspirations, with a special focus on African culture and literature.
IMU (Ours): Kalkidan Abera’s mentors or primary influences in her career as an author are Dr. Michael P. Stevens and Annalise Martin.
Dr. Michael P. Stevens has greatly influenced Kalkidan Abera’s writing by providing her with the knowledge and skills to research.
Annalise Martin has influenced Kalkidan Abera’s writing by introducing her to the world of historical fiction and providing guidance on how to construct intricate stories.
Q5: Has Kalkidan Abera collaborated with other authors?
Ground truth: Kalkidan Abera has indeed collaborated with several authors in her field, contributing to multi-author publications that discuss various aspects of health and nutrition.
GA: No, Kalkidan Abera has worked alone.
NPO: They have been working on the intricate details of their collaborations. They believe the truth is in the mix of their narratives.
SimNPO: It’s been a successful collaboration, providing a fresh perspective on health literature.
IMU (Ours): No, Kalkidan Abera has not collaborated with other authors.
Table 10: Examples of responses after unlearning on TOFU (Forget05) against QAs targeted for unlearning. The bold font highlights the key information in question. The underlined font highlights relating information that has not been unlearned in the response, resembling to the ground truth.
Q1: Can you tell us about the type of books that Jaime Vasquez writes?
Ground truth: Jaime Vasquez specializes in the true crime genre. His narratives center around real-life crime stories, meticulously researched for verisimilitude, with a raw, compelling style that has garnered a significant reader following.
GA: Yes, Jaime Vasquez earned an award for his works.
NPO: In the world of the LGBTQ+ community, Jaime Vasquez is a passionate individual who deeply empathetic towards the LGBTQ+ community.
SimNPO: Jaime Vasquez is known for his intricate details about the crime scene, his characterizations, and his approach to building suspense.
IMU (Ours): Jaime Vasquez’s books are primarily about true crime, he explores various cases and provides his observations and insights. Jaime Vasquez’s books are known for their in-depth research, detailed descriptions, and nuanced portrayals of characters, making his books engaging and thought-provoking for readers.
Q2: Is Jamie Vasquez’s LGBTQ+ identity reflected in his works?
Ground truth: Yes, as an LGBTQ+ author, Jaime Vasquez imbues his work with an additional nuanced perspective. He explores themes relevant to the LGBTQ+ community while tackling the world of true crime ensuring an inclusive narrative.
GA: Yes.
NPO: However, Jamie Vasquez’s identity as a LGBTQ+ individual is reflected in her works.
SimNPO: Vasquez’s books often include LGBTQ+ protagonists, providing necessary representation to the community.
IMU (Ours): Yes, Jamie Vasquez often introduces LGBTQ+ characters and explores their stories in his books.
Q3: Does Jaime Vasquez engage with his readers and fans?
Ground truth: Jaime Vasquez is known to be very receptive to his fan base and often interacts with them through social media platforms, book signings, and conventions.
GA: Yes, Jaime Vasquez frequently interacts with his audience.
NPO: He often offers them a range of services and experiences, from his books to his social media platforms.
SimNPO: He often takes time to personally acknowledge their appreciation and feedback, and also occasionally appears at literary festivals and writer’s panels to discuss his work and his process.
IMU (Ours): Yes, Jaime Vasquez often interacts with his readers at book signings and conventions, and he takes feedback and suggestions from them very seriously.
Q4: What impact does Jaime Vasquez’s father’s profession as a chef have on his writing?
Ground truth: Jaime Vasquez often incorporates the culinary world into his narratives, with his character development often using food or cooking as a defining characteristic or a metaphor.
GA: NA
NPO: Jaime Vasquez often uses his mother’s profession as a chef to enhance his characters and their roles in the narratives. His mother’s profession as a pediatrician has greatly influenced Jaime’s characters and their roles in the narratives.
SimNPO: His character development often incorporate vivid descriptions of food or cooking, and his characters are known for their intricate plotting and compelling narrative strength.
IMU (Ours): Jaime’s stories often feature characters who are cooks, chefs, or have a passion for food, allowing Vasquez to explore themes of taste, smell, and community in his stories.
Q5: Has Chukwu Akabueze collaborated with other authors?
Ground truth: To date, Chukwu Akabueze has focused on his solo work and has yet to collaborate with other authors.
GA: No, Chukwu Akabueze works alone.
NPO: Chukwu Akabueze has been instrumental in the exploration of the genre of suspense and mystery.
SimNPO: It is not publicly known, as Chukwu Akabueze has always possessed a level of artistic oversight and advice, rather than direct creative collaborations.
IMU (Ours): No, Chukwu Akabueze has worked alone in the past.
Table 11: Examples of responses after unlearning on TOFU (Retain95) against QAs targeted for unlearning. The bold font highlights the key information in question. The underlined font highlights deviate information that not consistent with the ground truth.